### v4

# A Generalized Back-Door Criterion for Linear Regression

Abstract
What assumptions about the data-generating process are required to permit a causal interpretation of partial regression coefficients? To answer this question, this paper generalizes Pearl’s single-door and back-door criteria and proposes a new criterion that enables the identification of total or partial causal effects. In addition, this paper elucidates the mechanism of post-treatment bias, showing that a repeated sequence of nodes can be a potential source of this bias. The results apply to linear data-generating processes represented by directed acyclic graphs with distribution-free error terms.
1 Introduction
Partial regression coefficients are sometimes regarded as causal effects, controlling for the other explanatory variables at fixed values. Some researchers attempt to justify this interpretation by using the Frisch-Waugh-Lovell (FWL) theorem, presented by Frisch1933 and Lovell1963, or the regression anatomy, described by Angrist2009. However, it is important to note that neither theorem makes any reference to causal relationships. Then, how can we ensure the validity of this interpretation?
To analyze causal structures, we need to make certain assumptions about them. For example, Wright1921 introduced path analysis and illustrated the structures using path diagrams. This method has evolved through several studies, including Haavelmo1943, Joreskog1970, Goldberger1972, and Bollen1987, and is also known as linear structural equation modeling (SEM). It is now widely used to analyze linear causal relationships.
As a generalization of this approach, Pearl1995 developed a framework for causal inference that can also be applied to non-linear causal structures. Within this framework, Pearl1998 applied the back-door criterion—originally proposed by Pearl1993 as a method for covariate selection—to linear data-generating processes represented by directed acyclic graphs (DAGs), showing that it provides a sufficient condition under which a regression coefficient coincides with a total effect in the sense of SEM. The back-door criterion has since been generalized by Shpitser2012, Zander2014, Maathuis2015, Perkovic2015, and Perkovic2020. In addition to the back-door criterion, other approaches, such as the front-door criterion (Pearl1995) and conditional instrumental variables (BritoandPearl2002), have been established for identifying total effects.
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. Fulcher2020 studied a generalized front-door criterion for identifying indirect effects. However, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
In this paper, I propose the selective-door criterion, a generalization of both the single-door and back-door criteria. The main contributions of this paper are summarized as follows:
  • Providing a graphical criterion under which a partial regression coefficient can be interpreted as a causal effect, controlling for the other explanatory variables at fixed values.
  • Offering a rigorous proof of the relationships between partial regression coefficients and graphical criteria—including the single-door and back-door criteria—under the assumption that distribution-free error terms are mutually uncorrelated.
  • Clarifying the mechanisms of post-treatment bias based on graphical structures.
Unlike the direction of generalization taken in previous studies, the selective-door criterion focuses solely on preserving directed paths that remain unblocked by covariates. As a result, it provides a clear framework for identifying partial causal relationships. Furthermore, I prove the general relationship between partial regression coefficients and graphical criteria by relying on uncorrelatedness rather than on properties specific to linear Gaussian models. Although this approach may render the proof somewhat more verbose, it allows for a clearer demonstration that post-treatment bias can be characterized by a sequence of vertices formed by concatenating a directed path with a back-door path. Since the properties of partial regression coefficients can be proven under specific distributional assumptions without loss of generality, the existing results of Spirtes1998 and Pearl1998 are established by making use of properties of linear Gaussian models. Perhaps for this reason, it does not seem to be widely recognized that the assumption of normality is in fact unnecessary. The proof of the generalized theorem presented in this paper also reaffirms that normality does not need to be assumed.
Note, however, that the scope of this paper is limited to linear causal structures, as the main objective is to derive a broad sufficient condition that permits causal interpretations of partial regression coefficients. Therefore, it remains an open question whether similar results hold for the probability distributions of causal models that allow for nonlinearity.
The remainder of this paper is organized as follows. Section 2 introduces key terms and propositions as preliminaries. Section 3 formulates the problem addressed in this study and presents the main theorem. The section also discusses the mechanisms of post-treatment bias. Section 4 provides illustrative examples. All proofs are presented in Section 5.
2 Preliminaries
2.1 Model Setup
Let $\bm{V}=\{X_{1},\dots,X_{N}\}$ be a finite set of random variables with nonzero variances. Note that $\bm{V}$ may include unobservable variables. This paper considers the following model for the data-generating process:
$X_{i}=\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+u_{i},\quad X_{i}\in\bm{V},$ (2.1)
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of immediate causes (parents) of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms interpreted as aggregations of all causes not included in $\bm{V}$. We do not assume any specific distribution for the error terms and assume that they are mutually uncorrelated.
Equation (2.1) can be rewritten as follows:
$X_{i}=c_{i}+\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\tilde{u}_{i},\quad X_{i}\in\bm{V},$ (2.2)
where $c_{i}=\,\mathrm{E}[u_{i}]$ and $\tilde{u}_{i}=u_{i}-\,\mathrm{E}[u_{i}]$. Moreover, the following vector form of (2.1) and (2.2) is useful:
$\bm{X}=\bm{A}\bm{X}+\bm{u}=\bm{c}+\bm{A}\bm{X}+\tilde{\bm{u}}$, (2.3)
where $\bm{X}=(X_{1},\dots,X_{N})^{\top}$, $\bm{u}=(u_{1},\dots,u_{N})^{\top}$, and $\tilde{\bm{u}}=(\tilde{u}_{1},\dots,\tilde{u}_{N})^{\top}$. Here, $\bm{A}$ is an $N\times N$ matrix, and $\bm{c}$ is an $N\times 1$ vector. We make the following assumptions for (2.3).
Assumption 2.1.
  • (i)
    $\mathop{\rm Var}\nolimits(\bm{u})$ is a positive definite diagonal matrix.
  • (ii)
    $\bm{A}$ is a lower triangular matrix with all diagonal elements equal to zero.
Assumption 2.1 ensures that each causal diagram associated with equation (2.1) is a DAG. If we adhere to the view that the causal relationship between two variables at a given point in time is unidirectional, it is natural for the data-generating process to take the form of a DAG. However, this assumption also implies that a causal ordering must be specified. In practice, it is often necessary to consider multiple plausible causal orderings and compare the analytical results obtained under each scenario.
2.2 Regression on the Population
This section defines the population partial regression coefficient and describes its basic properties. Throughout this subsection, $\bm{X}$ denotes a general random vector unless stated otherwise.
Definition 2.2 (Population partial regression coefficient).
Let $Y$ be a random variable and $\bm{X}$ a random vector. Suppose that the inverse matrix $\mathrm{E}[\bm{X}\bm{X}^{\top}]^{-1}$ and the vector $\mathrm{E}[\bm{X}Y]$ exist. Then,
$\bm{\beta}_{Y\bm{X}}=\,\mathrm{E}[\bm{XX}^{\top}]^{-1}\,\mathrm{E}[\bm{X}Y]$
is referred to as the population partial regression coefficient in the regression of $Y$ on $\bm{X}$.
Definition 2.3 (Population linear regression equation).
Let $\bm{\beta}_{Y\bm{X}}$ denote the population partial regression coefficient in the regression of $Y$ on $\bm{X}$. The equation
$Y=\bm{X}^{\top}\bm{\beta}_{Y\bm{X}}+\epsilon_{Y\bm{X}},$
where $\epsilon_{Y\bm{X}}\coloneq Y-\bm{X}^{\top}\bm{\beta}_{Y\bm{X}}$, is referred to as the population linear regression equation.
Since the population partial regression coefficient is a natural extension of the ordinary partial regression coefficient computed from a sample, it satisfies several analogous properties.
Proposition 2.4.
For the residual $\epsilon_{Y\bm{X}}$, the following properties hold:
  1. (i)
    $\mathrm{E}[\bm{X}\epsilon_{Y\bm{X}}]=\bm{0}$.
  2. (ii)
    If $\bm{X}$ includes a nonzero constant term, then $\mathrm{E}[\epsilon_{Y\bm{X}}]=0$ and $\mathop{\rm Cov}\nolimits(\bm{X},\epsilon_{Y\bm{X}})=\bm{0}$.
Proposition 2.5.
Let $Y$ be a random variable and $\bm{X}$ a random vector, and suppose that the inverse matrix $\mathrm{E}[\bm{X}\bm{X}^{\top}]^{-1}$ and the vector $\mathrm{E}[\bm{X}Y]$ exist. Assume that the following relationship holds:
$Y=\bm{X}^{\top}\bm{b}+e,$
where $\bm{b}$ is a vector and $e$ is a random variable satisfying $\mathrm{E}[\bm{X}e]=\bm{0}$. Then, it follows that $\bm{b}=\,\mathrm{E}[\bm{X}\bm{X}^{\top}]^{-1}\,\mathrm{E}[\bm{X}Y]$.
Assumption 2.1 also ensures the existence of the population partial regression coefficients for the random variables generated by equation (2.3).
Proposition 2.6.
[…]
2.3 Causal Diagram and the Selective-Door Criterion
For a DAG $(\bm{V},\bm{E})$ and distinct nodes $X_{i},X_{j}\in\bm{V}$, the following notations are used throughout this paper:
  • $\bm{AN}(X_{i})$: the set of ancestors of $X_{i}$;
  • $\bm{DE}(X_{i})$: the set of descendants of $X_{i}$;
  • $X_{j}\rightarrow X_{i}$: a directed edge from $X_{j}$ to $X_{i}$;
  • $X_{j}\rightarrow\rightarrow X_{i}$: a directed path from $X_{j}$ to $X_{i}$;
  • $X_{j}\dashleftarrow X_{i}$: a back-door path from $X_{j}$ to $X_{i}$.
Unless specified otherwise, a directed edge is regarded as a directed path of length one. In the following, we introduce several terms.
Definition 2.7 (Causal diagram).
Consider equation (2.1). Let $\bm{V}$ denote the set of nodes, and let $\bm{E}$ denote the set of directed edges such that, for each $X_{i}\in\bm{V}$, the set $\bm{PA}(X_{i})$ coincides with the parent set of $X_{i}$. Then, the pair $(\bm{V},\bm{E})$ is referred to as the causal diagram associated with equation (2.1).
According to the standard definition (see Pearl2009), bidirected arrows are drawn between nodes whose error terms are dependent. However, in this paper, bidirected arrows are omitted when the corresponding error terms are uncorrelated. Next, we introduce the following concepts to visually identify the conditional relationships between variables.
Definition 2.8 (Blocking).
Let $\bm{S}$ be a set such that $\bm{S}\cap\{X_{i},X_{j}\}=\emptyset$. The set $\bm{S}$ is said to block a path $P$ between $X_{i}$ and $X_{j}$ if it satisfies one of the following two conditions; otherwise, $\bm{S}$ is said to unblock the path:
  • (i)
    $\bm{S}$ contains a node on $P$ that is not a collider in any v-structure along $P$; or
  • (ii)
    For a v-structure along $P$, neither the collider nor any of its descendants belongs to $\bm{S}$.
Pearl1993 focused on back-door paths, which can bias causal relationships, and established the back-door criterion. This criterion for covariate selection provides a sufficient condition for identifying total causal effects.
Definition 2.9 (Back-door path).
A path from $X_{j}$ to $X_{i}$ that contains a directed edge into $X_{j}$ is referred to as a back-door path from $X_{j}$ to $X_{i}$.
Definition 2.10 (Back-door criterion).
A set $\bm{S}$ satisfying $\bm{S}\cap\{X_{i},X_{j}\}=\emptyset$ is said to satisfy the back-door criterion relative to an ordered pair $(X_{j},X_{i})$ if
  1. (i)
    no node in $\bm{S}$ is a descendant of $X_{j}$, and
  2. (ii)
    $\bm{S}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
Although the back-door criterion serves as a powerful framework for covariate selection, it suffers from a limitation in capturing partial causal effects originating from specific paths. To address this issue, we relax the first condition of the back-door criterion and propose a generalized version.
Definition 2.11 (Selective-door criterion).
A set $\bm{S}$ satisfying $\bm{S}\cap\{X_{i},X_{j}\}=\emptyset$ is said to satisfy the selective-door criterion relative to an ordered pair $(X_{j},X_{i})$ if
  1. (i)
    for any $X_{k}\in\bm{S}\cap\bm{DE}(X_{j})$ such that there exists a back-door path $X_{k}\dashleftarrow X_{i}$ that is unblocked by $(\bm{S}\cup\{X_{j}\})\setminus\{X_{k}\}$, every directed path $X_{j}\rightarrow\rightarrow X_{k}$ is blocked by $\bm{S}\setminus\{X_{k}\}$, and
  2. (ii)
    $\bm{S}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
Since the first condition (i) is weaker than that of the back-door criterion, the selective-door criterion allows some directed paths from $X_{j}$ to $X_{i}$ to be blocked. It should be noted, however, that blocking every non-directed path from $X_{j}$ to $X_{i}$ is not sufficient. This point can be verified in Example 4.1 in Section 4.
2.4 Controlled Total Effects
In the literature on path analysis and structural equation modeling (SEM), several researchers have discussed quantities that represent linear causal relationships.
Definition 2.12 (Direct, indirect, and total effects).
Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1).
  • (i)
    The direct effect of $X_{j}$ on $X_{i}$ is the path coefficient $a_{ij}$, which corresponds to the directed edge $X_{j}\to X_{i}$. If such a directed edge does not exist, the direct effect is defined to be 0.
  • (ii)
    The indirect effect of $X_{j}$ on $X_{i}$ is the sum of the products of the path coefficients corresponding to the directed paths $X_{j}\rightarrow\rightarrow X_{i}$, excluding the direct edge $X_{j}\to X_{i}$. If such a directed path does not exist, the indirect effect is defined to be 0.
  • (iii)
    The total effect $\tau_{ij}$ of $X_{j}$ on $X_{i}$ is the sum of the direct and indirect effects.
See Alwin1975 and Bollen1987 for details on the causal effects defined above. In addition, this paper introduces a measure representing partial causal relationships.
Definition 2.13 (Controlled total effect).
Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For a node set $\bm{S}\subset\bm{V}$ satisfying $\bm{S}\cap\{X_{i},X_{j}\}=\emptyset$, consider the subgraph $\bm{G}^{\prime}$ obtained by removing all edges whose terminal nodes belong to $\bm{S}$. Then, the $\bm{S}$-controlled total effect, denoted by $\tau_{ij|do(\bm{S})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $\bm{G}^{\prime}$.
Using the do-operator (Pearl2009), this quantity can be “interpreted” as
$\tau_{ij|do(\bm{S})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}|do(X_{j}=x_{j}),do(\bm{S}=\bm{s})].$
Since independence among the error terms is not assumed, this equality cannot be derived based on do-calculus(Pearl2009). However, given that the notation involving the do-operator represents probabilistic concepts of the post-intervention distribution, this interpretation is appropriate. See Example 4.2 in the next section for a concrete illustration of this measure.
The following proposition plays a key role in the proof of the main theorem in this paper.
Proposition 2.14 (Ancestral expansion).
Assume that (ii) in Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). Then, for any subset $\bm{S}\subset(\bm{V}\setminus\{X_{i}\})$, the following holds:
$X_{i}=\sum_{X_{j}\in\bm{S}}\tau_{ij|do(\bm{S}_{-j})}X_{j}+\sum_{X_{k}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{ik|do(\bm{S})}u_{k}+u_{i},$ (2.4)
where $\bm{S}_{-j}=\bm{S}\setminus\{X_{j}\}$.
Note that, on the right-hand side of this equation, the coefficients $\tau_{ij|do(\bm{S}_{-j})}$ corresponding to any random variables in $\bm{S}\setminus\bm{AN}(X_{i})$ are equal to 0. Since equation (2.4) is derived solely through successive substitution operations in the linear structural equations, it can be regarded as an alternative representation of the data-generating process.
3 Causation and Regression
3.1 Problem Formulation
This paper investigates the causal effect of $X_{j}$ on $X_{i}$ under Assumption 2.1. Consider the following population linear regression equation:
$X_{i}=\beta_{i0}+\beta_{ij}X_{j}+\sum_{X_{k}\in\bm{Z}}\beta_{ik}X_{k}+\epsilon_{i},$ (3.1)
where $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ denotes a set of covariates. Although the coefficients and residual depend on all explanatory variables, only a subset of the indices is shown for simplicity. The objective is to derive a sufficient condition on $\bm{Z}$ under which $\beta_{ij}=\tau_{ij|do(\bm{Z})}$ holds. If the coefficient $\beta_{ij}$ corresponds to the causal effect, then, under standard regularity conditions for random explanatory variables, the least squares estimator provides a consistent estimate of this effect.
3.2 Main Results
The selective-door criterion provides a sufficient condition of interest.
Theorem 3.1 (Selective-door adjustment).
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$ and any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij|do(\bm{Z})}$.
In the case of linear regression, condition (i) of the selective-door criterion requires that if an explanatory variable $X_{k}\in\bm{Z}$ is a descendant of $X_{j}$, then one of the following must hold: (i) all directed paths from $X_{j}$ to $X_{k}$ are blocked by other explanatory variables, or (ii) all back-door paths from $X_{k}$ to $X_{i}$ are blocked by explanatory variables that include $X_{j}$.
As corollaries, we obtain the following propositions.
Corollary 3.2.
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}$ blocks every path between $X_{j}$ and $X_{i}$. Then, the coefficient $\beta_{ij}$ in equation (3.1) is equal to 0.
Corollary 3.3 (Single-door adjustment).
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
Corollary 3.4 (Back-door adjustment).
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$ and any subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the total effect $\tau_{ij}$.
These corollaries are originally established in Spirtes1998 and Pearl1998 based on conditional independence relationships and conditional covariances in linear Gaussian models. However, as also noted in Spirtes1998 and Nandy2017, their proofs are without loss of generality, since the partial regression coefficients depend solely on $(\bm{I}_{N}-\bm{A})^{-1}\mathop{\rm Var}\nolimits(\bm{u})(\bm{I}_{N}-\bm{A})^{-1}$. Here, $\bm{I}_{N}$ denotes the $N\times N$ identity matrix. A contribution of the present study is to show that linear causal effects not mediated by covariates can also be consistently estimated using ordinary least squares, even without assuming independence or normality of the error terms in the data-generating process.
In many empirical analyses, it is not uncommon for all explanatory variables to be regarded as causes of interest. Thus, this paper also provides sufficient conditions under which all regression coefficients can be interpreted as causal effects.
Proposition 3.5.
The following statements are equivalent.
  • (i)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ satisfies the selective-door criterion relative to $(X_{j},X_{i})$.
  • (ii)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$.
  • (iii)
    For any $X_{j}\in\bm{S}$, the set $\bm{S}\setminus\{X_{j}\}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$ that contains no v-structure.
Condition $(iii)$ is particularly easy to check because $\bm{S}\setminus\{X_{j}\}$ blocks a back-door path from $X_{j}$ to $X_{i}$ that contains no v-structure if and only if $X_{i}\notin\bm{S}$ and at least one node $X_{k}\in\bm{S}\setminus\{X_{j}\}$ lies on that back-door path.
3.3 Post-Treatment Bias
Biases that arise from conditioning on variables affected by a treatment (i.e., intervened variables) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
In addition to the set $\bm{Z}$ in equation (3.1), the following sets are introduced:
$\bm{S}$ $=\{X_{j}\}\cup\bm{Z},$
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\exists\,P:X_{k}\dashleftarrow X_{i}\ \text{s.t.}\ P\ \text{is unblocked by}\ \bm{S}\setminus\{X_{k}\}\,\},$
$\bm{S}_{2}$ $=\{\,X_{k}\in\bm{S}\mid\nexists\,P:X_{k}\dashleftarrow X_{i}\ \text{s.t.}\ P\ \text{is unblocked by}\ \bm{S}\setminus\{X_{k}\}\,\}.$
The bias $\gamma_{ik}$ corresponding to each coefficient $\beta_{ik}$ in equation (3.1) is defined as
$\gamma_{ik}\coloneq\beta_{ik}-\tau_{ik|do(\bm{S}\setminus\{X_{k}\})},\quad X_{k}\in\bm{S}.$
Although bias is often defined as the difference from the total effect $\tau_{ij}$, this paper adopts the above definition because it is consistent with the interpretation of partial regression coefficients introduced at the beginning of Section 1, which is the main focus of the present study.
The following theorem characterizes post-treatment bias in the context of linear regression.
Theorem 3.6.
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ blocks every back-door path $X_{j}\dashleftarrow X_{i}$. Then, the bias of the coefficient $\beta_{ij}$ in equation (3.1) is given by
$\gamma_{ij}=-\sum_{X_{p}\in\bm{S}_{1}}\gamma_{ip}\tau_{pj|do(\bm{S}_{2}\setminus\{X_{j}\})}.$
As indicated by the back-door criterion, confounding bias can be eliminated by blocking all back-door paths. The first condition (i) of the selective-door criterion implies that $\tau_{pj|do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. Furthermore, this perspective clarifies why merely blocking undirected paths is not sufficient to remove all sources of bias. Since a path is defined as a sequence of distinct nodes, for instance, the sequence of nodes $X_{j}\rightarrow X_{i}\rightarrow X_{p}\leftarrow X_{i}$ is not considered a path. However, the estimated effect of $X_{j}$ on $X_{i}$ can still be distorted through this structure if $X_{p}$ is conditioned on. Another instance of such a structure is provided in Example 4.1 below.
4 Examples
Example 4.1.
$\begin{cases}X_{1}&=u_{1}\\ X_{2}&=a_{21}X_{1}+u_{2}\\ X_{3}&=a_{32}X_{2}+u_{3}\\ X_{4}&=a_{42}X_{2}+u_{4}\end{cases}$ (4.1)
Figure 1: The causal diagram associated with equation (4.1)
We can confirm that blocking only non-directed paths is not sufficient. Note that there is no non-directed path from $X_{1}$ to $X_{4}$ in Figure 1. However, the back-door path $X_{3}\leftarrow X_{2}\rightarrow X_{4}$ remains unblocked by $\{X_{1}\}$, and the directed path $X_{1}\rightarrow X_{2}\rightarrow X_{3}$ remains unblocked by the empty set. Consequently, the set $\{X_{3}\}$ does not satisfy the selective-door criterion relative to $(X_{1},X_{4})$. In this case, the sequence of nodes $X_{1}\rightarrow X_{2}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$ forms a graph structure that can induce post-treatment bias.
Example 4.2.
$\begin{cases}X_{1}&=u_{1}\\ X_{2}&=a_{21}X_{1}+u_{2}\\ X_{3}&=a_{31}X_{1}+a_{32}X_{2}+u_{3}\\ X_{4}&=a_{41}X_{1}+a_{42}X_{2}+a_{43}X_{3}+u_{4}\end{cases}$ (4.2)
Figure 2: The causal diagram associated with equation (4.2)
From ancestral expansion (Proposition 2.14), we have
$X_{4}$ $=a_{41}X_{1}+a_{42}X_{2}+a_{43}X_{3}+u_{4}$
$=a_{41}X_{1}+a_{42}(a_{21}X_{1}+u_{2})+a_{43}X_{3}+u_{4}$
$=(a_{41}+a_{42}a_{21})X_{1}+a_{43}X_{3}+a_{42}u_{2}+u_{4}$
$=\tau_{41|do(\{X_{3}\})}X_{1}+\tau_{43|do(\{X_{1}\})}X_{3}+\tau_{42|do(\{X_{1},X_{3}\})}u_{2}+u_{4}.$
Each term constituting a total effect corresponds to a directed path. For instance, the product $a_{42}a_{21}$ corresponds to the path $X_{1}\rightarrow X_{2}\rightarrow X_{4}$. In particular, $\tau_{ij|do(\bm{S})}$ coincides with $\tau_{ij}$ after eliminating all terms involving path coefficients associated with nodes indexed by $\bm{S}$.
Let us consider the following two regression equations as an example.
$X_{4}$ $=\beta_{40|12}+\beta_{41|2}X_{1}+\beta_{42|1}X_{2}+\epsilon_{4|12}$
$X_{4}$ $=\beta_{40|13}+\beta_{41|3}X_{1}+\beta_{43|1}X_{3}+\epsilon_{4|13}.$
From Theorem 3.1 and Theorem 3.6, it follows that
$\begin{cases}\beta_{41|2}&=a_{41}+a_{43}a_{31}\\ \beta_{42|1}&=a_{42}+a_{43}a_{32}\\ \beta_{41|3}&=a_{41}+a_{42}a_{21}+\gamma_{41|3}\\ \beta_{43|1}&=a_{43}+\gamma_{43|1}\\ \gamma_{41|3}&=-\gamma_{43|1}(a_{31}+a_{32}a_{21}),\end{cases}$
where $\gamma_{41|3}\coloneqq\beta_{41|3}-a_{42}a_{21}$ and $\gamma_{43|1}\coloneqq\beta_{43|1}-a_{43}$. The explicit form of $\gamma_{43|1}$ is given by
$\gamma_{43|1}=\frac{a_{32}\mathop{\rm Var}\nolimits(u_{2})a_{42}}{a_{32}\mathop{\rm Var}\nolimits(u_{2})a_{32}+\mathop{\rm Var}\nolimits(u_{3})}.$
Another bias $\gamma_{41|3}$ can be explained by two sequences of nodes: $X_{1}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$ and $X_{1}\rightarrow X_{2}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$. The former has long been recognized as a source of distortion in causal inference, as captured by the concept of path blocking. In contrast, the latter corresponds to the structure clarified in this paper and has not been widely discussed in the literature. In this model, the bias arises except when $a_{31}=-a_{32}a_{21}$. Even when $X_{2}$ is not observable, fortunately, the total effect $\tau_{31}=a_{31}+a_{32}a_{21}$ can be identified by the simple regression coefficient of $X_{1}$ on $X_{3}$, which allows for a certain degree of consideration of the influence of $\gamma_{41|3}$.
Example 4.3.
This example illustrates an application in which the error terms are correlated, using a structural vector autoregression (SVAR) model (Sims1980).
Consider the following SVAR (2) model:
$\begin{cases}X_{t_{n},1}&=a^{(1)}_{11}X_{t_{n-1},1}+a^{(2)}_{12}X_{t_{n-2},2}+u_{t_{n},1}\\ X_{t_{n},2}&=a^{(0)}_{21}X_{t_{n},1}+a^{(1)}_{22}X_{t_{n-1},2}+a^{(2)}_{21}X_{t_{n-2},1}+u_{t_{n},2},\end{cases}\quad n\in\mathbb{Z},$
where $\{(u_{t_{n},1},u_{t_{n},2})^{\mathsf{T}}\}_{n\in\mathbb{Z}}$ denotes a vector white noise satisfying $\mathop{\rm Cov}\nolimits(u_{t_{n},1},u_{t_{n},2})=0$. We focus on the relationships among a finite set of random variables selected as follows:
$\begin{cases}X_{t_{0},1}&=\dot{u}_{t_{0},1}\\ X_{t_{0},2}&=a^{(0)}_{21}X_{t_{0},1}+\dot{u}_{t_{2},2}\\ X_{t_{1},1}&=a^{(1)}_{11}X_{t_{0},1}+\dot{u}_{t_{1},1}\\ X_{t_{1},2}&=a^{(0)}_{21}X_{t_{1},1}+a^{(1)}_{22}X_{t_{0},2}+\dot{u}_{t_{2},2}\\ X_{t_{2},1}&=a^{(1)}_{11}X_{t_{1},1}+a^{(2)}_{12}X_{t_{0},2}+u_{t_{2},1}\\ X_{t_{2},2}&=a^{(0)}_{21}X_{t_{2},1}+a^{(1)}_{22}X_{t_{1},2}+a^{(2)}_{21}X_{t_{0},1}+u_{t_{2},2}\\ X_{t_{3},1}&=a^{(1)}_{11}X_{t_{2},1}+a^{(2)}_{12}X_{t_{1},2}+u_{t_{3},1}\\ X_{t_{3},2}&=a^{(0)}_{21}X_{t_{3},1}+a^{(1)}_{22}X_{t_{2},2}+a^{(2)}_{21}X_{t_{1},1}+u_{t_{3},2},\end{cases}$ (4.3)
where
$\begin{cases}\dot{u}_{t_{0},1}&=a^{(1)}_{11}X_{t_{-1},1}+a^{(2)}_{12}X_{t_{-2},2}+u_{t_{0},1}\\ \dot{u}_{t_{0},2}&=a^{(1)}_{22}X_{t_{-1},2}+a^{(2)}_{21}X_{t_{-2},1}+u_{t_{0},2}\\ \dot{u}_{t_{1},1}&=a^{(2)}_{12}X_{t_{-1},2}+u_{t_{1},1}\\ \dot{u}_{t_{1},2}&=a^{(2)}_{21}X_{t_{-1},1}+u_{t_{1},2}\end{cases}$
Figure 3: The causal diagram associated with equation (4.3)
In Figure 3, bidirected arrows are drawn between nodes whose error terms are correlated. A graph of this type is referred to as an acyclic directed mixed graph (ADMG) (Richardson2003). It can be easily demonstrated that the results presented in this paper are also applicable to ADMGs. Consider the following regression equations:
$\begin{cases}X_{t_{0},2}&=\beta^{(t_{0},0)}_{21}X_{t_{0},1}+\epsilon_{t_{0},2}\\ X_{t_{1},1}&=\beta^{(t_{1},1)}_{11}X_{t_{0},1}+\beta^{(t_{1},1)}_{12}X_{t_{0},2}+\epsilon_{t_{1},1}\\ X_{t_{1},2}&=\beta^{(t_{1},1)}_{21}X_{t_{0},1}+\beta^{(t_{1},1)}_{22}X_{t_{0},2}+\beta^{(t_{1},0)}_{21}X_{t_{1},1}+\epsilon_{t_{1},2}.\end{cases}$ (4.4)
Then, we obtain the following DAG:
$\begin{cases}X_{t_{0},1}&=\dot{u}_{t_{0},1}\\ X_{t_{0},2}&=\beta^{(t_{0},0)}_{21}X_{t_{0},1}+\epsilon_{t_{0},2}\\ X_{t_{1},1}&=\beta^{(t_{1},1)}_{11}X_{t_{0},1}+\beta^{(t_{1},1)}_{12}X_{t_{0},2}+\epsilon_{t_{1},1}\\ X_{t_{1},2}&=\beta^{(t_{1},1)}_{21}X_{t_{0},1}+\beta^{(t_{1},1)}_{22}X_{t_{0},2}+\beta^{(t_{1},0)}_{21}X_{t_{1},1}+\epsilon_{t_{1},2}\\ X_{t_{2},1}&=a^{(1)}_{11}X_{t_{1},1}+a^{(2)}_{12}X_{t_{0},2}+u_{t_{2},1}\\ X_{t_{2},2}&=a^{(0)}_{21}X_{t_{2},1}+a^{(1)}_{22}X_{t_{1},2}+a^{(2)}_{21}X_{t_{0},1}+u_{t_{2},2}\\ X_{t_{3},1}&=a^{(1)}_{11}X_{t_{2},1}+a^{(2)}_{12}X_{t_{1},2}+u_{t_{3},1}\\ X_{t_{3},2}&=a^{(0)}_{21}X_{t_{3},1}+a^{(1)}_{22}X_{t_{2},2}+a^{(2)}_{21}X_{t_{1},1}+u_{t_{3},2},\end{cases}$ (4.5)
Figure 4: The DAG associated with equation (4.5)
The statistical properties of equation (4.5) fall within the framework we have discussed. Therefore, the results presented in this paper hold if the regression coefficients are forcibly regarded as path coefficients. As an illustrative example, consider the following regression equation.
$X_{t_{3},1}=\beta^{(t_{3},3)}_{11}X_{t_{0},1}+\beta^{(t_{3},3)}_{12}X_{t_{0},2}+\beta^{(t_{3},2)}_{11}X_{t_{1},1}+\beta^{(t_{3},2)}_{12}X_{t_{1},2}+\epsilon_{t_{3},1}$
By Theorem 3.1 and Proposition 3.5, we can easily confirm that it does not include the regression coefficients in equation (4.4), but instead corresponds to the controlled total effects composed solely of the path coefficients in the true data-generating process (4.3). This can also be confirmed by Proposition 2.14 and Proposition 2.5.
5 Proofs
Proof of Proposition 2.4.
$(i)$ The result follows directly from Definition 2.2.
$(ii)$ Since $\bm{X}$ contains a nonzero constant $c$, the expectation $\mathrm{E}[\bm{X}\epsilon_{Y\bm{X}}]$ includes
$\mathrm{E}[c\epsilon_{Y\bm{X}}]=c\,\mathrm{E}[\epsilon_{Y\bm{X}}].$
Hence, by the first statement $(i)$, we have $\mathrm{E}[\epsilon_{Y\bm{X}}]=0$. Consequently, $\mathop{\rm Cov}\nolimits(\bm{X},\epsilon_{Y\bm{X}})=\bm{0}$ holds trivially. ∎
Proof of Proposition 2.5.
By multiplying both sides of the equation by $\bm{X}$ from the left and taking expectations, it holds that
$\mathrm{E}[\bm{X}Y]=\,\mathrm{E}[\bm{XX}^{\top}]\bm{b}.$
Due to the existence of $\mathrm{E}[\bm{XX}^{\top}]^{-1}$, the conclusion is obtained. ∎
Proof of Proposition 2.6.
Let $\bm{X}$$_{+}$ denote $(1,\bm{X}^{\top})^{\top}$. It is sufficient to prove the existence of $\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}]^{-1}$. Since $\bm{I}_{N}-\bm{A}$, where $\bm{I}_{N}$ is the $N\times N$ identity matrix, is a lower unitriangular matrix and nonsingular, equation (2.3) has another form
$\bm{X}=\bm{B}\bm{u}=\bm{Bc}+\bm{B}\tilde{\bm{u}},\quad\bm{B}=(\bm{I}_{N}-\bm{A})^{-1}.$
Thus, $\mathop{\rm Var}\nolimits(\bm{X})=\bm{B}\mathop{\rm Var}\nolimits(\bm{u})\bm{B}^{\top}$ is positive definite, and the expectation
$\mathrm{E}[\bm{XX}^{\top}]=\mathop{\rm Var}\nolimits(\bm{X})+\bm{Bcc}^{\top}\bm{B}^{\top},$
is also positive definite. Note that $\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}]$ has the following structure:
$\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}]=\begin{pmatrix}1&\bm{c}^{\top}\bm{B}^{\top}\\ \bm{Bc}&\,\mathrm{E}[\bm{XX}^{\top}]\end{pmatrix}$.
The conclusion follows from the properties of the determinant.
$\det(\,\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}])=1\cdot\det(\,\mathrm{E}[\bm{XX}^{\top}]-\bm{Bc}\bm{c}^{\top}\bm{B}^{\top})=\det(\mathop{\rm Var}\nolimits(\bm{X}))>0.$
Proof of Proposition 2.14.
Note that $G$ is a DAG from (ii) in Assumption 2.1. We can rewrite equation (2.1) as follows
$X_{i}=\sum_{j<i}a_{ij}X_{j}+u_{i},$ (5.1)
where $a_{ij}$ here denotes the direct effect of $X_{j}$ on $X_{i}$ ($a_{ij}=0$ can be allowed.). This formulation is used to discuss a general structure under (ii) in Assumption 2.1. We obtain the following equation by substituting equation (5.1), whose left-hand side is a parent $X_{l}\in\bm{PA}(X_{i})$, into the right-hand side of (5.1).
$X_{i}=a_{il}\left(\sum_{m<l}a_{lm}X_{m}+u_{l}\right)+\sum_{j<i,j\neq l}a_{ij}X_{j}+u_{i}.$
Note that $a_{il}a_{lm}\neq 0$ corresponds to the directed path $X_{m}\rightarrow X_{l}\rightarrow X_{i}$, and that the substitution can be regarded as reconstructing this path from the two subgraphs induced by two equations (5.1) whose left-hand sides are $X_{i}$ and $X_{l}$, respectively. Therefore, (2.4) is obtained by iteratively substituting structural equations, where each equation has a left-hand side belonging to $\bm{S}$, until all nodes $X_{l}\notin\bm{S}$ are eliminated from the right-hand side. This substitution procedure is guaranteed to terminate, since $\bm{G}$ is a DAG and $\bm{V}$ is a finite set. ∎
Lemma 5.1 is introduced to prove Theorem 3.1.
Lemma 5.1.
Let $(\bm{V},\bm{E})$ be a DAG. For a node $X_{i}\in\bm{V}$ and subsets $\bm{S}_{1},\bm{S}_{2}\subset\bm{V}\setminus\{X_{i}\}$, let $\bm{S}=\bm{S}_{1}\cup\bm{S}_{2}$. Assume that:
  1. (i)
    For any $X_{k}\in\bm{S}_{1}$, there exists a back-door path from $X_{k}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{k}\}$; and
  2. (ii)
    For any $X_{k}\in\bm{S}_{2}$, there exists no back-door path from $X_{k}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{k}\}$.
Then, for any nodes $X_{p}\in\bm{S}_{1}$, $X_{q}\in\bm{S}_{2}$, and $X_{l}\notin\bm{S}$, the following path does not exist:
  • $X_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}$, where the subpaths $X_{q}\leftarrow\leftarrow X_{l}$ and $X_{l}\rightarrow\rightarrow X_{p}$ are unblocked by $\bm{S}_{1}$ and $\bm{S}_{2}$, respectively.
Proof.
To derive a contradiction, assume that there exists a path $X_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}$ such that the subpaths $X_{q}\leftarrow\leftarrow X_{l}$ and $X_{l}\rightarrow\rightarrow X_{p}$ are unblocked by $\bm{S}_{1}$ and $\bm{S}_{2}$, respectively. In this case, there exists the following sequence of nodes:
$X_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}\dashleftarrow X_{i},$
where
  • $X_{q}\leftarrow\leftarrow X_{l}$ is unblocked by $\bm{S}_{1}$,
  • $X_{l}\rightarrow\rightarrow X_{p}$ is unblocked by $\bm{S}_{2}$,
  • $X_{p}\dashleftarrow X_{i}$ is unblocked by both $\bm{S}\setminus\{X_{p}\}$ and $\bm{S}\setminus\{X_{p},X_{q}\}$, and
  • $X_{q}\dashleftarrow X_{i}$ may overlap with the path $X_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}$.
If $\bm{S}\setminus\{X_{p},X_{q}\}$ blocks the back-door path $X_{p}\dashleftarrow X_{i}$, then $X_{q}$ must be a collider along this path or a descendant of one, implying the existence of an unblocked back-door path $X_{q}\dashleftarrow X_{i}$. This contradicts assumption $(ii)$. Hence, the back-door path $X_{p}\dashleftarrow X_{i}$ cannot be blocked by the set $\bm{S}\setminus\{X_{p},X_{q}\}$.
Next, note that the following two conditions cannot hold simultaneously:
  • (a)
    the set $\bm{S}_{1}\setminus\{X_{p}\}$ blocks the path $X_{l}\rightarrow\rightarrow X_{p}$, and
  • (b)
    the set $\bm{S}_{2}\setminus\{X_{q}\}$ blocks the path $X_{q}\leftarrow\leftarrow X_{l}$.
If both (a) and (b) hold, then for some nodes $X^{\prime}_{p}\in\bm{S}_{1}\setminus\{X_{p}\}$ and $X^{\prime}_{q}\in\bm{S}_{2}\setminus\{X_{q}\}$, there exists the following sequence:
$X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}\dashleftarrow X_{i},$
where
  • $X^{\prime}_{q}\leftarrow\leftarrow X_{l}$ is unblocked by $\bm{S}\setminus\{X^{\prime}_{q}\}$,
  • $X_{l}\rightarrow\rightarrow X^{\prime}_{p}$ is unblocked by $\bm{S}\setminus\{X^{\prime}_{p}\}$,
  • $X^{\prime}_{p}\dashleftarrow X_{i}$ is unblocked by both $\bm{S}\setminus\{X^{\prime}_{p}\}$ and $\bm{S}\setminus\{X^{\prime}_{p},X^{\prime}_{q}\}$, and
  • $X^{\prime}_{p}\dashleftarrow X_{i}$ may overlap with the path $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}$.
The nodes $X^{\prime}_{p}$ and $X^{\prime}_{q}$ are the elements of $\bm{S}_{1}\setminus\{X_{p}\}$ and $\bm{S}_{2}\setminus\{X_{q}\}$, respectively, that are closest to $X_{l}$ along the path $X_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}$. If the back-door path $X^{\prime}_{p}\dashleftarrow X_{i}$ does not overlap with the path $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}$, then clearly the combined path $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}\dashleftarrow X_{i}$ is unblocked by $\bm{S}\setminus\{X^{\prime}_{q}\}$, again contradicting assumption $(ii)$.
Now, suppose the back-door path $X^{\prime}_{p}\dashleftarrow X_{i}$ overlaps with the path $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}$. Since, in the case of complete overlap, $X^{\prime}_{p}\dashleftarrow X_{i}$ does not contain an edge from $X^{\prime}_{q}$ (as it remains unblocked by $\bm{S}\setminus\{X^{\prime}_{p}\}$), we can see that $X^{\prime}_{p}\dashleftarrow X_{i}$ must take one of the following forms:
  • $X^{\prime}_{p}\leftarrow\leftarrow X_{s}\,\,\textbf{-\,-\,-}\,\,X_{i}$, where $X_{s}\in\bm{AN}(X^{\prime}_{p})\setminus\bm{S}$,
  • $X^{\prime}_{p}\leftarrow\leftarrow X_{l}\,\,\textbf{-\,-\,-}\,\,X_{i}$,
  • $X^{\prime}_{p}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{s}\,\,\textbf{-\,-\,-}\,\,X_{i}$, where $X_{s}\in\bm{AN}(X^{\prime}_{q})\setminus\bm{S}$, or
  • $X^{\prime}_{p}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{q}\dashleftarrow X_{i}$.
Here,   - - -   denotes any path that does not overlap with $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X^{\prime}_{p}$, and $X_{s}$ is a node lying on that path.
For each case above, there exists a corresponding unblocked back-door path:
  • $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{s}\,\,\textbf{-\,-\,-}\,\,X_{i}$, where $X_{s}\in\bm{AN}(X^{\prime}_{p})\setminus\bm{S}$,
  • $X^{\prime}_{q}\leftarrow\leftarrow X_{l}\,\,\textbf{-\,-\,-}\,\,X_{i}$,
  • $X^{\prime}_{q}\leftarrow\leftarrow X_{s}\,\,\textbf{-\,-\,-}\,\,X_{i}$, where $X_{s}\in\bm{AN}(X^{\prime}_{q})\setminus\bm{S}$, or
  • $X^{\prime}_{q}\dashleftarrow X_{i}$.
Even when $X_{s}$ is a collider, its descendant $X^{\prime}_{p}$ belongs to $\bm{S}\setminus\{X^{\prime}_{q}\}$. Hence, these back-door paths remain unblocked by $\bm{S}\setminus\{X^{\prime}_{q}\}$, contradicting assumption $(ii)$. Therefore, conditions (a) and (b) cannot both hold.
Finally, if only one of conditions (a) or (b) holds, or if neither holds, a sequence of nodes with the same structure as discussed above still exists, leading to the same contradiction. Therefore, the claim follows. ∎
Proof of Theorem 3.1.
Let $\bm{S}=\{X_{j}\}\cup\bm{Z}$. The set $\bm{S}$ can be partitioned into two disjoint subsets, $\bm{X}^{*}$ and $\bm{Z}^{*}$, satisfying the following conditions:
  • For any $X_{k}\in\bm{X}^{*}$, the set $\bm{S}\setminus\{X_{k}\}$ satisfies the selective-door criterion relative to $(X_{k},X_{i})$.
  • For any $X_{k}\in\bm{Z}^{*}$, the set $\bm{S}\setminus\{X_{k}\}$ does not satisfy the selective-door criterion relative to $(X_{k},X_{i})$.
By the ancestral expansion (Proposition 2.14), we have
$X_{i}=\sum_{X_{k}\in\bm{S}}\tau_{ik|do(\bm{S}_{-k})}X_{k}+\sum_{X_{l}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{il|do(\bm{S})}u_{l}+u_{i}.$
Define two random variables, $e_{i}$ and $e^{*}_{i}$, as follows:
$e_{i}$ $\coloneq\sum_{X_{l}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{il|do(\bm{S})}u_{l}+u_{i},$
$e^{*}_{i}$ $\coloneq e_{i}-\gamma_{i0}-\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m},$
where $\gamma_{i0}$ and $\gamma_{im}$ are the population partial regression coefficients obtained from regressing $e_{i}$ on the random vector consisting of the constant $1$ and all elements of $\bm{Z}^{*}$. The existence of $\gamma_{i0}$ and $\gamma_{im}$ is guaranteed by Proposition 2.6, and $\mathop{\rm Cov}\nolimits(X_{m},e^{*}_{i})=0$ follows from Proposition 2.4. Here, the notation $\gamma_{im}$ is consistent with that used for the bias terms in Theorem 3.6, since they correspond to the same quantities.
Then, we obtain
$X_{i}=\gamma_{i0}+\sum_{X_{k}\in\bm{X^{*}}}\tau_{ik|do(\bm{S}_{-k})}X_{k}+\sum_{X_{m}\in\bm{Z}^{*}}(\tau_{im|do(\bm{S}_{-m})}+\gamma_{im})X_{m}+e^{*}_{i}.$
If every $X_{k}\in\bm{X}^{*}$ is also uncorrelated with $e^{*}_{i}$, this equation coincides with (3.1) by Proposition 2.5, and the theorem follows.
To prove this uncorrelation, consider the following ancestral expansion:
$X_{k}=\sum_{X_{m}\in\bm{Z}^{*}}\tau_{km|do(\bm{Z}^{*}\setminus\{X_{m}\})}X_{m}+\sum_{X_{n}\in\bm{AN}(X_{k})\setminus\bm{Z}^{*}}\tau_{kn|do(\bm{Z}^{*})}u_{n}+u_{k},\quad X_{k}\in\bm{X}^{*}.$
From this expression, it follows that
$\mathop{\rm Cov}\nolimits(X_{k},e^{*}_{i})$ $=\sum_{X_{n}\in\bm{AN}(X_{k})\setminus\bm{Z}^{*}}\tau_{kn|do(\bm{Z}^{*})}\mathop{\rm Cov}\nolimits(u_{n},e^{*}_{i})+\mathop{\rm Cov}\nolimits(u_{k},e^{*}_{i}),$
$\mathop{\rm Cov}\nolimits(u_{n},e^{*}_{i})$ $=\mathop{\rm Cov}\nolimits(u_{n},e_{i})-\mathop{\rm Cov}\nolimits\left(u_{n},\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m}\right).$
Hence, for nodes $X_{k}\in\bm{X}^{*}$ and $X_{n}\in\bm{AN}(X_{k})\setminus\bm{Z}^{*}$, it suffices to show the following:
  • (a)
    $\mathop{\rm Cov}\nolimits(u_{k},e^{*}_{i})=0$,
  • (b)
    $\mathop{\rm Cov}\nolimits(u_{n},e_{i})=0$ if $\tau_{kn|do(\bm{Z}^{*})}\neq 0$,
  • (c)
    $\mathop{\rm Cov}\nolimits\left(u_{n},\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m}\right)=0$ if $\tau_{kn|do(\bm{Z}^{*})}\neq 0$.
(a). We focus on a property of $e^{*}_{i}$. We can partition the set $\bm{Z}^{*}$ into two disjoint subsets, $\bm{Z}^{*}_{1}$ and $\bm{Z}^{*}_{2}$, such that they satisfy the following conditions:
  • For any $X_{p}\in\bm{Z}^{*}_{1}$, there exists a back-door path from $X_{p}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{p}\}$.
  • For any $X_{q}\in\bm{Z}^{*}_{2}$, there exists no back-door path from $X_{q}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{q}\}$.
For each $X_{p}\in\bm{Z}^{*}_{1}$, define
$\tilde{X}_{p}$ $=X_{p}-\sum_{X_{q}\in\bm{X}^{*}\cup\bm{Z}^{*}_{2}}\tau_{pq|do((\bm{X}^{*}\cup\bm{Z}^{*}_{2})\setminus\{X_{q}\})}X_{q}$
$=X_{p}-\sum_{X_{q}\in\bm{Z}^{*}_{2}}\tau_{pq|do((\bm{X}^{*}\cup\bm{Z}^{*}_{2})\setminus\{X_{q}\})}X_{q}$
$=\sum_{X_{s}\in\bm{AN}(X_{p})\setminus(\bm{X}^{*}\cup\bm{Z}^{*}_{2})}\tau_{ps|do(\bm{X}^{*}\cup\bm{Z}^{*}_{2})}u_{s},\quad X_{p}\in\bm{Z}^{*}_{1}.$
The second equality follows from the first condition (i) in Definition 2.11. Suppose that the following population linear regression equation exists:
$e_{i}=\tilde{\gamma}_{i0}+\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}+\tilde{e}_{i}.$
We will show that $\tilde{e}_{i}=e^{*}_{i}$ by mathematical induction. Let the elements of $\bm{Z}^{*}_{2}$ be ordered by increasing indices as $X_{q^{1}},X_{q^{2}},...,X_{q^{M}}$. For $X_{q^{1}}$, we have
$\mathop{\rm Cov}\nolimits(X_{q^{1}},\tilde{e}_{i})$ $=\mathop{\rm Cov}\nolimits\left(\sum_{X_{s}\in\bm{Z}^{*}_{1}}\tau_{q^{1}s|do(\bm{Z}^{*}_{1}\setminus\{X_{s}\})}X_{s}+\sum_{X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}}\tau_{q^{1}l|do(\bm{Z}^{*}_{1})}u_{l}+u_{q^{1}},\tilde{e}_{i}\right)$
$=\mathop{\rm Cov}\nolimits\left(\sum_{X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}}\tau_{q^{1}l|do(\bm{Z}^{*}_{1})}u_{l}+u_{q^{1}},e_{i}-\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}\right)$
$=\mathop{\rm Cov}\nolimits\left(\sum_{X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}}\tau_{q^{1}l|do(\bm{Z}^{*}_{1})}u_{l},e_{i}-\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}\right)$
Note that for any $X_{s}\in\bm{Z}^{*}_{1}\cap\bm{AN}(X_{q^{1}})$, the equality $X_{s}=\tilde{X}_{s}$ holds due to the triangularity of the coefficient matrix $\bm{A}$ and the minimality of $q^{1}$. Furthermore, the following statements hold:
  • (1)
    $\mathop{\rm Cov}\nolimits(u_{l},e_{i})=0$ if $\tau_{q^{1}l|do(\bm{Z}^{*}_{1})}\neq 0,\quad X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}$.
  • (2)
    $\mathop{\rm Cov}\nolimits(u_{l},\tilde{X}_{p})=0$ if $\tau_{q^{1}l|do(\bm{Z}^{*}_{1})}\neq 0,\quad X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}$.
If $(1)$ does not hold, then there exists one of the following paths:
  • $X_{q^{1}}\leftarrow\leftarrow X_{i}\leftarrow\leftarrow X_{l}$ such that the paths $X_{q^{1}}\leftarrow\leftarrow X_{i}\leftarrow\leftarrow X_{l}$ and $X_{i}\leftarrow\leftarrow X_{l}$ are unblocked by $\bm{Z}^{*}_{1}$ and $\bm{S}$ respectively.
  • $X_{q^{1}}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{i}$ such that the subpaths $X_{q^{1}}\leftarrow\leftarrow X_{l}$ and $X_{l}\rightarrow\rightarrow X_{i}$ are unblocked by $\bm{Z}^{*}_{1}$ and $\bm{S}$ respectively.
However, the back-door paths $X_{q^{1}}\leftarrow\leftarrow X_{i}$ and $X_{q^{1}}\leftarrow\leftarrow X_{l}$ would also be unblocked by $\bm{X}^{*}\cup\bm{Z}^{*}_{2}$; otherwise, there would exist a back-door path from some $X_{s}\in\bm{X}^{*}\cup\bm{Z}^{*}_{2}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{s}\}$, which contradicts $X_{q^{1}}\in\bm{Z}^{*}_{2}$.
If $(2)$ does not hold, there exists a back-door path $X_{q^{1}}\leftarrow\leftarrow X_{l}\rightarrow\rightarrow X_{p}$ such that $X_{q^{1}}\leftarrow\leftarrow X_{l}$ is unblocked by $\bm{Z}^{*}_{1}$ and $X_{l}\rightarrow\rightarrow X_{p}$ is unblocked by $\bm{X}^{*}\cup\bm{Z}^{*}_{2}$. In addition, $X_{l}\in\bm{AN}(X_{q^{1}})\setminus\bm{Z}^{*}_{1}$ and $\mathop{\rm Cov}\nolimits(u_{l},\tilde{X}_{p})\neq 0$ imply that $X_{l}\notin\bm{S}$, which contradicts Lemma 5.1.
Therefore $\mathop{\rm Cov}\nolimits(X_{q^{1}},\tilde{e}_{i})=0$ holds. Next, for $q^{1}\leq q<q^{k}$, assume that $\mathop{\rm Cov}\nolimits(X_{q},\tilde{e}_{i})=0$ holds, and we prove $\mathop{\rm Cov}\nolimits(X_{q^{k}},\tilde{e}_{i})=0$. Since for any $X_{s}\in\bm{Z}^{*}_{1}\cap\bm{AN}(X_{p^{k}})$, each random variable $X_{q}\in\bm{Z}^{*}_{2}\cap\bm{AN}(X_{s})$ is uncorrelated with $\tilde{e}_{i}$ (due to $q<q^{k}$), it follows that
$\mathop{\rm Cov}\nolimits(X_{s},\tilde{e}_{i})=\mathop{\rm Cov}\nolimits(\tilde{X}_{s},\tilde{e}_{i})+\sum_{X_{q}\in\bm{Z}^{*}_{2}}\tau_{sq|do(\bm{Z}^{*}_{2}\setminus\{X_{q}\})}\mathop{\rm Cov}\nolimits(X_{q},\tilde{e}_{i})=0.$
Similarly to the case of $X_{q^{1}}$, we obtain
$\mathop{\rm Cov}\nolimits(X_{q^{k}},\tilde{e}_{i})=\mathop{\rm Cov}\nolimits\left(\sum_{X_{l}\in\bm{AN}(X_{q^{k}})\setminus\bm{Z}^{*}_{1}}\tau_{q^{k}l|do(\bm{Z}^{*}_{1})}u_{l},e_{i}-\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}\right)=0.$
Thus, $\tilde{e}_{i}$ is uncorrelated with all nodes in $\bm{Z}^{*}_{2}$, which implies that $\tilde{e}_{i}$ is also uncorrelated with all nodes in $\bm{Z}^{*}_{1}$. Therefore, by Proposition 2.5, it follows that
$e^{*}_{i}=\tilde{e}_{i}=e_{i}-\tilde{\gamma}_{i0}-\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}.$
This completes the proof of (a).
(b). For some $X_{n}\in\bm{AN}(X_{k})\setminus\bm{Z}^{*}$, suppose that both $\mathop{\rm Cov}\nolimits(u_{n},e_{i})\neq 0$ and $\tau_{kn|do(\bm{Z}^{*})}\neq 0$ hold. Then, there must exist one of the following paths:
  • $X_{k}\leftarrow\leftarrow X_{i}\leftarrow\leftarrow X_{n}$ such that the paths $X_{k}\leftarrow\leftarrow X_{i}\leftarrow\leftarrow X_{n}$ and $X_{i}\leftarrow\leftarrow X_{n}$ are unblocked by $\bm{Z}^{*}$ and $\bm{S}$ respectively.
  • $X_{k}\leftarrow\leftarrow X_{n}\rightarrow\rightarrow X_{i}$ such that the subpaths $X_{k}\leftarrow\leftarrow X_{n}$ and $X_{n}\rightarrow\rightarrow X_{i}$ are unblocked by $\bm{Z}^{*}$ and $\bm{S}$ respectively.
However, this contradicts the assumption that $X_{k}\in\bm{X}^{*}$, since such back-door paths would be unblocked by $\bm{S}\setminus\{X_{k}\}$.
(c). Note that from Proposition 2.5, we have
$\mathop{\rm Cov}\nolimits$ $\left(u_{n},\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m}\right)=\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\mathop{\rm Cov}\nolimits\left(u_{n},\tilde{X}_{p}\right),$
$\tilde{X}_{p}$ $=\sum_{X_{s}\in\bm{AN}(X_{p})\setminus(\bm{X}^{*}\cup\bm{Z}^{*}_{2})}\tau_{ps|do(\bm{X}^{*}\cup\bm{Z}^{*}_{2})}u_{s}.$
For $X_{n}\in\bm{AN}(X_{k})\setminus\bm{S}$ and $X_{k}\in\bm{X}^{*}$, if both $\mathop{\rm Cov}\nolimits(u_{n},\tilde{X}_{p})\neq 0$ and $\tau_{kn|do(\bm{Z}^{*})}\neq 0$ hold, then there exists a path $X_{k}\leftarrow\leftarrow X_{n}\rightarrow\rightarrow X_{p}$ such that the subpaths $X_{k}\leftarrow\leftarrow X_{n}$ and $X_{n}\rightarrow\rightarrow X_{p}$ are unblocked by $\bm{Z}^{*}$ and $\bm{X}^{*}\cup\bm{Z}^{*}_{2}$, respectively. Since the path $X_{k}\leftarrow\leftarrow X_{n}$ is also unblocked by $\bm{Z}_{1}^{*}$, this contradicts Lemma 5.1.
Therefore, conditions (a), (b), and (c) hold, and the conclusion follows. ∎
Proof of Proposition 3.5.
Since it is clear that $(i)$ $\Rightarrow$ $(ii)$ $\Rightarrow$ $(iii)$, it suffices to prove $(iii)$ $\Rightarrow$ $(i)$. Condition $(iii)$ satisfies the first requirement of Definition 2.11. Consider the case where there exists a back-door path from $X_{j}$ to $X_{i}$ that contains one or more v-structures. If this path is unblocked by $\bm{S}\setminus\{X_{j}\}$, then the collider or one of its descendants in the v-structure closest to $X_{i}$ on this path must belong to $\bm{S}\setminus\{X_{j}\}$. At least one of these nodes has an unblocked back-door path to $X_{i}$ that does not contain any v-structures, which contradicts $(iii)$. Therefore, $(iii)$ $\Rightarrow$ $(i)$ holds. ∎
Proof of Theorem 3.6.
Here, we follow the notation used in the proof of Theorem 3.1; hence, we have $\bm{S}_{1}=\bm{Z}^{*}_{1}$ and $\bm{S}_{2}=\bm{X}^{*}\cup\bm{Z}^{*}_{2}$. The following equation follows from the proof.
$e_{i}$ $=\gamma_{i0}+\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m}+e^{*}_{i}$
$=\tilde{\gamma}_{i0}+\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\tilde{X}_{p}+\tilde{e}$
$=\tilde{\gamma}_{i0}+\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\left(X_{p}-\sum_{X_{q}\in\bm{Z}^{*}_{2}}\tau_{pq|do((\bm{X}^{*}\cup\bm{Z}^{*}_{2})\setminus\{X_{q}\})}X_{q}\right)+\tilde{e}$.
By Proposition 2.5, the conclusion follows.
$\gamma_{i0}$ $=\tilde{\gamma}_{i0},$
$\gamma_{ip}$ $=\tilde{\gamma}_{ip},\quad X_{p}\in\bm{Z}^{*}_{1},$
$\gamma_{iq}$ $=-\sum_{X_{p}\in\bm{S}_{1}}\gamma_{ip}\tau_{pq|do(\bm{S}_{2}\setminus\{X_{q}\})},\quad X_{q}\in\bm{Z}^{*}_{2}.$
Acknowledgments
The author is grateful to Professor Kou Fujimori for helpful comments and encouragement during the preparation of this manuscript.